OpenStudy (anonymous):

A wildlife manager determines that there are approximately 200 deer in a certain state park. a. The population is growing at a rate of 7% per year. How many deer will in the park after 4 years? b. If the carrying capacity of this park is 350 deer, how long will it take for the deer population to reach carrying capacity?

OpenStudy (anonymous):

</3

OpenStudy (anonymous):

7% in decimal form is .07 1st year 200+200(.07)=200(1+.07) nth year 200(1.07)^n a) plug in 4 for n above b)200(1.07)^n=350 (solve for n)